Jake throws two dice, one green and one red. What is the probability that the green die is greater than the red die number?
Answer by stanbon(75887)

Jake throws two dice, one green and one red. What is the probability that the green die is greater than the red die number?
----
The mean result for the red die is (1/6)(1+2+3+4+5+6) = 21/6 = 3.5
---
P(green > 3.5) = 3/6 = 1/2
===============================
Cheers,
Stan H.
=============